### v2

# On Focusing Statistical Power for Searches and Measurements in Particle Physics

Abstract
Particle physics experiments rely on the (generalised) likelihood ratio test (LRT) for searches and measurements, which consist of composite hypothesis tests. However, this test is not guaranteed to be optimal, as the Neyman-Pearson lemma pertains only to simple hypothesis tests. Any choice of test statistic thus implicitly determines how statistical power varies across the parameter space. An improvement in the core statistical testing methodology for general settings with composite tests would have widespread ramifications across experiments. We discuss an alternate test statistic that provides the data analyzer an ability to focus the power of the test on physics-motivated regions of the parameter space. We demonstrate the improvement from this technique compared to the LRT on a Higgs $\rightarrow\tau\tau$ dataset simulated by the ATLAS experiment and a dark matter dataset inspired by the LZ experiment. We also employ machine learning to efficiently perform the Neyman construction, which is essential to ensure statistically valid confidence intervals.
I Introduction
A major theme in high-energy physics is the search for new phenomena and the desire to precisely measure known physics parameters. The design of such experiments can be expensive, and it can take decades to achieve the required sensitivity to weak or rare signals in an ocean of uninteresting backgrounds. A search in this context concludes when the existence of a particle has been confidently established after analyzing enough collisions, usually by rejecting the default hypothesis that the observed excess in a given sample would occur by chance, as with the Higgs boson discovery [ATLAS:2012yve, CMS:2012qbp]. Alternatively, the search may conclude by confidently ruling out the existence of a certain class of particles, as with several supersymmetry and dark matter searches [LZ:2024zvo, Kwon:2025fgb]. Similarly, a measurement is reported in the form of a confidence interval on a key property of the particle — a smaller interval means a more precise measurement.
Experimental particle physicists have almost universally relied on the generalized likelihood ratio test (LRT) for such studies [cousins_lectures_2024, cranmer_practical_2015]. The LRT is optimal for simple-vs-simple hypothesis tests (Neyman-Pearson’s lemma, [neyman_ix_1933]), and also has convenient asymptotic properties [wilks_large-sample_1938, Cowan:2010js].
However, searches and measurements are typically performed as composite hypothesis tests. For instance, the measurement of a signal strength parameter for a potential new phenomenon or the mass of a known particle generally involves testing a default conjecture, such as $H_{0}:\mu=\mu_{0}$, against a composite alternative hypothesis, such as $H_{1}:\mu\neq\mu_{0}$, which encompasses a range of possible parameter values rather than a single one. In these cases, the Neyman–Pearson lemma no longer applies: the most powerful test may vary depending on the specific value $\mu_{1}$ within the alternative. In fact, for general settings, there is no single test that is simultaneously most powerful across all parameter values in the composite alternative. Effectively, this means that any choice of test statistic is implicitly a choice for how the statistical power of the test varies over the parameter space; no test will be the most powerful everywhere.
In this Letter, we present an alternative to LRT that allows the data analyst to focus the power of a test on specific regions of the parameter space — without assuming simple hypotheses or large sample sizes, while maintaining validity of confidence intervals. Our proposed focused test statistic (FTS) compares the null hypothesis to a weighted average of all alternative hypotheses, where the weighting is defined by a physics-motivated “focus function” that increases the sensitivity of the test to the regions of the parameter space of most interest to physicists. We show that the FTS outperforms the LRT statistic for a collider physics example and a dark matter search example.
For our two case studies, we achieve significant gains using FTS in regimes with small sample sizes and low signal fractions, which is precisely where the most interesting physics can be found. Our method constructs confidence intervals via a Neyman construction, which is commonly implemented using a computationally expensive Feldman-Cousins technique [Feldman:1997qc] in particle physics. We also demonstrate a machine learning approach to rapidly perform Neyman construction [pmlr-v119-dalmasso20a, dalmasso_likelihood-free_2024, masserano_simulator-based_2023] that provides valid confidence intervals for complex particle physics datasets. This technique applies to either test statistic and does not require large sample sizes, thereby enabling the construction of valid confidence intervals in a range of dark matter, neutrino or collider experiments where local asymptotics cannot be assumed [cranmer_practical_2015, NOvA:2017abs, ATLAS:2024jry, LZ:2025xkj].
As a drop-in replacement for the LRT statistic, FTS can be applied to any existing particle physics analysis and holds the potential to improve sensitivity in key measurements across experiments, such as neutrino oscillation parameters, Higgs self-coupling, Higgs width, effective field theory parameters and also searches, such as for dark matter.
This Letter is organized as follows: Section II introduces the statistical setup. Section III compares the two test statistics for a Higgs measurement and a dark matter search problem. Section IV discusses our main results.
II Statistical setup
II.1 How do we construct confidence intervals from tests?
Searches and measurements in particle physics are typically performed as hypothesis tests on the parameter $\mu$ of the likelihood $p(\mathcal{D}|\mu)$ of datum $\mathcal{D}$ implied by the physics processes under study for some parameter space $\Theta$ (See Appendix A for a detailed description of the experimental setup). For example, the generalized LRT rejects the null hypothesis $H_{0}:\mu=\mu_{0}$ (default conjecture on, e.g., the strength $\mu$ of a signal physics process) in favor of the alternative hypothesis $H_{1}:\mu\neq\mu_{0}$ when the generalized likelihood ratio statistic (LRS)
$T(\mathcal{D};\mu_{0})=-2\log\left(\frac{p(\mathcal{D}|\mu_{0})}{\sup_{\mu\in\Theta}p(\mathcal{D}|\mu)}\right)$ (1)
is large or above some “critical value” $C_{\mu_{0}}$, which is chosen so that the proportion of wrongly rejected events (“false positives”) do not exceed some pre-set significance level $\alpha$. These hypothesis tests can then also be converted into an interval estimate of $\mu$ via the Neyman construction [neyman_outline_1937], where a $(1-\alpha)100$% confidence interval $I(\mathcal{D})$ of $\mu$ is defined by the set of all parameter values that passed the level-$\alpha$ tests:
$I(\mathcal{D})=\{\mu_{0}\in\Theta\mid T(\mathcal{D};\mu_{0})<C_{\mu_{0}}\}.$ (2)
See Figure 1 (top) for a depiction of the construction. The statistical efficiency or “power” of the procedure (which determines how tight the intervals are) depends on the intricacies of the likelihood and the choice of the test statistic $T$. The computational efficiency of the procedure depends on how fast we can compute the test statistic $T(\mathcal{D};\mu_{0})$ and the critical values $C_{\mu_{0}}$ as a function of $\mu_{0}$. In practice, $\Theta$ is a discrete but high resolution grid of values for $\mu_{0}$. Furthermore, to avoid boundary effects, we replace $\Theta$ in the denominator of Equation 1 by an extended parameter space $\widetilde{\Theta}$, which is defined as the parameter space $\Theta$ padded with one unit at both ends; see Appendix A.1 for details.
II.2 How do we cheaply compute critical values?
Data analyses in collider physics often rely on asymptotic properties (such as Wilks’ theorem [wilks_large-sample_1938, Cowan:2010js]) to cheaply estimate LRS confidence intervals, but asymptotics do not hold in small sample regimes. In this Letter, we propose a machine learning (ML) approach to quickly estimate how $C_{\mu_{0}}$ might vary with $\mu_{0}$; see Algorithm 1 in Appendix B for details of a procedure that performs a quantile regression of $T$ on $\mu_{0}$. An ML-based approach allows the data analyst to (i) take advantage of alternative test statistics (such as the proposed FTS defined below in Eq. 3), (ii) explore challenging small-sample regimes, and (iii) extend HEP analyses to higher-dimensional parameter spaces (such as in effective field theory) or to unbinned analyses, where a Monte Carlo (MC) approach (e.g., [cranmer_practical_2015, Sec. 5.4]) is not computationally feasible.
II.3 What test statistic can increase power in physics-motivated parameter regions?
LRS is optimal for simple-vs-simple tests but is not guaranteed to be optimal for more general settings with high power against all alternatives. Hence, we propose an alternative test statistic (referred to as a focused test statistic; FTS) of the form
$T_{f}(\mathcal{D};\mu_{0})=-2\log\left(\frac{p(\mathcal{D}|\mu_{0})}{\int_{\Theta}p(\mathcal{D}|\mu)f(\mu)d\mu}\right),$ (3)
which allows the data analyst to focus the power of a test or confidence procedure to specific regions of the parameter space, as defined by the focus or weight function $f$. In Bayesian testing, this statistic is known as a Bayes factor. Whereas the idea of using the Bayes factor as a frequentist test statistic is not new [good_bayesnon-bayes_1992, dalmasso_likelihood-free_2024, Fowlie03082023], this work is the first to demonstrate that a test based on the Bayes factor has the potential to improve upon likelihood ratio tests in various particle physics settings. For simplicity, we will in our experiments use a Gaussian focus function $f(\mu)$ with mean $m$ and standard deviation $s$, truncated so that $f(\mu)=0$ for $\mu\notin\Theta$. This choice may be guided by prior knowledge of $\mu_{0}$, for example, from an independent previous measurement or from theoretical considerations. The construction of an FTS confidence interval follows the procedure discussed above; see Figure 1 (bottom).
III Experiments and Results
We highlight two applications – a measurement of the coupling of Higgs bosons to tau leptons and a search for the existence of weakly interacting massive particles (WIMPs). The likelihoods are computed from bins of a histogram according to the counting experiment convention as described in Appendix A. We compare LRS confidence intervals with FTS where we use truncated Gaussian focus functions of various widths (both narrow and wide choices), to demonstrate the robustness of our qualitative conclusions to these choices. To compute critical values, we adopt the procedure described in Section II.1.
Figure 1: Construction of confidence intervals. (Top) To construct a $(1-\alpha)100\%$ confidence interval using the LRS statistic (yellow; Eq. 1), we estimate the critical values $C_{\mu_{0}}$ (grey) via quantile regression and retain those $\mu_{0}$ values for which the statistic falls below the critical value. Shown are the critical values for $68\%$ (dark grey) and $95\%$ (light grey) confidence levels, with corresponding intervals displayed below the figure (dark and light yellow, respectively). (Bottom) The FTS statistic (blue; Eq. 3) for a focus function centered at $m=1$ and with width $s=1.2$ (dashed red) is compared against the critical values. Intervals at $68\%$ and $95\%$ confidence level are shown (dark and light blue, respectively). This figure corresponds to the Higgs measurement example using the Higgs mass observable, with the test statistics are evaluated on an Asimov data set [Cowan:2010js]
III.1 Measurement of a Higgs coupling via its signal strength
The goal of this case study is to measure the coupling of the Higgs boson to tau leptons [Aad:2015vsa]. This unknown quantity has a one-to-one correspondence with the signal strength in the Standard Model (SM) of particle physics and can therefore be posed as a signal strength measurement. The default hypothesis for the true signal strength $\mu^{*}=1$ corresponds to the SM prediction. For this study, we let $\Theta=[0,10]$ to avoid right endpoint truncation bias in interval estimation, and we set the center of the focus function at $m=1$, the SM prediction. We define the wide focus width to be $s=2.4$, corresponding to the $2\sigma$ expected sensitivity of a prior measurement [ATLAS:2012gfw]; a narrow focus width of $s=1.2$ is also considered.
III.1.1 Dataset
The dataset was simulated by the ATLAS experiment for the 2014 Higgs boson machine learning challenge (HiggsML) [HiggsMLOpenData, pmlr-v42-cowa14]. It is comprised of a mixture of events obtained via several Higgs decay and background processes. For further details on the challenge, refer to [pmlr-v42-cowa14].
In order to show the robustness of the analysis to our experimental design choices, we examine two natural choices of observables to build template histograms for the counting experiment setup. The first observable is the reconstructed mass of the Higgs boson candidate as estimated via phase space integration; the second is the invariant mass of the hadronic tau and final-state lepton (“visible mass” for short). Both observables are known to discriminate between signal and background with the reconstructed mass expected to provide stronger constraints.
III.1.2 Results
When measuring the Higgs coupling with the reconstructed Higgs mass, FTS with either a wide ($s=2.4$) or narrow ($s=1.2$) focus gives 68% confidence intervals [^1] The 68% and 95% confidence intervals discussed throughout this Letter correspond to $1\sigma$ (68.27%) and $2\sigma$ (95.45%) intervals, which is the convention in high-energy physics. that have shorter median length than LRS. Figure 2(a) shows the median interval lengths as they vary with the true value of $\mu^{*}$ while the focus function remains unchanged, with center fixed at $m=1$. The improvement is robust to a misspecification of the focus function (meaning that the center of the focus function $m$ differs from the true value of the parameter $\mu^{*}$) and the confidence intervals remain valid by construction. When $\mu^{*}=1$, the FTS intervals for a wide focus are about 13% shorter than those obtained via LRS; see Table 1. With a narrow focus function, they are 21% shorter than LRS intervals. Using the visible mass observable to constrain the parameter of interest, FTS shows a decrease in median length by 13% with wide focus and 22% with narrow focus. Table 1 shows that the absolute reduction in median interval length of FTS over LRS is roughly the same at the 68% and 95% levels. However, since 95% intervals are longer than 68% intervals, this also implies that the proportional gain is smaller at 95% than at 68% (about a 6% reduction for FTS-wide using the mass observable, 11% for FTS-narrow; and a 6% reduction for FTS-wide using visible mass, 10% for FTS-narrow).
Estimation of critical values for both the LRS and FTS is accelerated by using quantile regression. With a low simulation budget of about 9,0009{,}000 pseudo-experiments (PEs), which for MC corresponds to 30 PEs per point times $300$ grid points, the mean-squared error (MSE) is improved by a factor of about 6 for LRS and 3 for FTS. See Figure 3 for an illustration of the relative speed-up.
III.2 Search for WIMPs
Several experiments are dedicated to the search for dark matter candidates [Bozorgnia:2024pwk]. We design a data analysis scenario inspired by the search for WIMPs at the LZ experiment [LZ:2019sgr]. For our study, we construct a simplified dataset that mimics the signal and background densities, as well as the sample size, of the recent LZ dark matter search [LZ:2024zvo]. The default hypothesis is that these dark matter candidates do not exist, corresponding to $\mu^{*}=0$; exclusion of this hypothesis would constitute a discovery. For this reason, our focus functions are centered at $m=0$. In this study, we let $\Theta=[0,40]$ to avoid upper endpoint truncation bias in interval estimation; the focus function is thus truncated below at 0.
III.2.1 Dataset
We produce synthetic data mimicking the full dataset that passes selection cuts (as seen in [LZ:2024zvo, Fig. 3]) relating densities of background and 40 GeV WIMP models (see Appendix A.2 for details). For consistency with the previous example, we use two-dimensional histograms for density estimation, although these can readily be replaced with an analytical model for the density, when they are available in dark matter experiments [Aalbers:2020iej]. The resulting 2D template histograms for the signal and background distributions are given in two observables, $S_{1}$ and $\log_{10}(S_{2})$, representing the light from prompt vacuum ultraviolet scintillation and delayed electroluminescence in the detector (see Fig. 4 in Appendix A.2). The simulated data are normalized such that when $\mu^{*}=1$, there are 1,2001{,}200 background events per $1$ signal event. The NEST simulator [brodsky_nestcollaborationnest_2019] was used to generate the template histograms.
III.2.2 Results
In Figure 2b we show median confidence interval length in the LZ-inspired experiment as a function of the true signal strength. We find that FTS improves on LRS across the range of $0\leq\mu^{*}\leq 15$ with both wide ($s=6.0$) and narrow ($s=3.0$) focus functions. Table 1 (bottom two rows) shows that FTS with wide focus provides a median length $22\%$ shorter than LRS for a 68% confidence level in the setting where $\mu^{*}=0$. With narrow focus, the FTS intervals are $35\%$ shorter. As a result, we would expect a true negative search to conclude sooner with our method relative to the LRS-based method, on average. Furthermore, FTS maintains an advantage over LRS when $\mu^{*}\neq 0$. When $\mu^{*}=1$, the median $68\%$ confidence interval lengths for FTS are still $25\%$ shorter than LRS intervals when using wide focus and $34\%$ with narrow focus. If WIMPs exist, our method would find evidence for them sooner by excluding the default hypothesis using less data. See Appendix C for a discussion on setting upper and lower limits. As in the Higgs case, the absolute gains in median interval length are similar for 68% and 95% confidence levels. For $\mu^{*}=0$, there is 17% reduction for FTS-wide, 19% for FTS-narrow; and for $\mu^{*}=1$, a 16% reduction for FTS-wide and 16% for FTS-narrow.
Figure 2: Lengths of confidence intervals. (a, left) Results for the Higgs measurement with wide focus ($s=2.4$). The thick lines represent the median interval length with the uncertainty bands denoting 25% and 75% quantiles of the length distribution. FTS (blue) yields 13% shorter intervals than LRS (orange) near the focus center ($m=1$). (a, right) For a narrow focus ($s=1.2$), FTS intervals are about 25% shorter. That advantage is maintained even at a modest distance from the center, e.g. near $\mu^{*}=2$. (b, left-right) For the LZ-inspired search, the intervals are shorter across the domain of our search. Near $\mu^{*}=0$, the FTS intervals are about 35% shorter with wide focus than LRS, and 22% shorter with narrow focus.
Setting Test statistic
Experiment $\mu^{*}$ LRS FTS-wide FTS-narrow
Higgs (mass) 1.0 1.08 (2.01) 0.94 (1.89) 0.85 (1.79)
Higgs (vis. mass) 1.0 1.26 (2.31) 1.10 (2.17) 0.98 (2.08)
LZ-inspired 0.0 5.99 (13.87) 4.68 (11.47) 3.89 (11.29)
LZ-inspired 1.0 7.08 (15.18) 5.29 (12.82) 4.68 (12.75)
Table 1: Median confidence interval length (in units of $\mu$) at the 68% (95%) confidence level. For the Higgs examples, the focus center is $m=1.0$; the focus widths are $s=2.4$ for FTS-wide, and $s=1.2$ for FTS-narrow. For LZ-inspired examples, the focus center is $m=0$; $s=6.0$ for FTS-wide, and $s=3.0$ for FTS-narrow. In each setting, the best result (the shortest median interval length) is bold-faced.
Figure 3: Critical values. (a) We compare the mean-squared error (MSE) in the critical value estimates obtained via MC (red) and quantile regression (QR, grey) for LRS on a grid of 300300 evaluation points. For a low simulation budget of 9,0009{,}000 pseudo-experiments (PEs), MC estimates show high MSE whereas QR estimates (using the same number of PEs) are comparable to MC estimates with $1.35$ million PEs. The right column confirms that low-budget QR yields accurate estimates across the parameter space, unlike low-budget MC. (b) For FTS, QR is again more efficient. With 9,0009{,}000 PEs, QR matches the performance of MC with about 30,00030{,}000 PEs.
IV Discussion and Conclusion
In this Letter, we propose the focus test statistic (FTS; Equation 3) as a drop-in replacement for the likelihood ratio statistic (LRS; Equation 1), which is widely used in high-energy physics for both parameter measurements and searches for new phenomena. These analyses rely on composite hypotheses, where the Neyman–Pearson lemma does not guarantee optimality of the LRS across the parameter space. In two proof-of-concept studies, a Higgs measurement at a collider experiment and a WIMP search in a dark matter experiment, we demonstrate that FTS can outperform LRS by focusing the power of the test in physics-motivated regions of the parameter space. FTS with a narrow, well-specified focus function yields 21% smaller 68% confidence intervals than LRS for the Higgs measurement (when $\mu^{*}$ is at the SM point) and 35% smaller intervals for the dark matter search (when there is no WIMP signal). For the dark matter search, we also find that FTS yields tighter upper bounds on the signal strength when the signal does not exist as well as higher lower bounds when it does, for a fixed amount of collected data; see Appendix C. In our study, we take the variability between PEs into account and report the median interval length over PEs, together with the 25th and 75th percentiles of the length distribution.
The FTS framework gives the scientist control over the regions of maximum sensitivity, whether chosen after optimization studies on independent simulations, informed by independent previous measurements (as in our Higgs example) or by theory, without affecting the validity of the confidence intervals. As a theory-driven example, if an angular parameter is being measured, it may be sensible to concentrate power within $0,2\pi]$ using, e.g., a von Mises (circular normal) distribution. The focus function allows the analyst to leverage prior knowledge about reasonable parameter regions within a fully frequentist analysis. However, the data under analysis should not be used to determine the focus. For example, one can explore the impact of different focus functions on the sensitivity using simulated samples over a range of true values of the parameter (see Figure 2) to provide guidelines, if the simulations did not use the observed data. The general rule is that a narrow focus leads to larger gains in power around the center of mass of the focus function, but one then pays a price (as compared to a wider focus) in terms of decreased power if the focus center of mass differs from the true value of the parameter $\mu^{*}$, as illustrated by the top row of Figure 2.
As an improvement to the core statistical methodology of the field, the FTS can enhance discovery potential and physics sensitivities across many experiments. Our case studies suggest that significant gains can be achieved when the signal-to-noise ratio is low and sample size is small, such as for dark matter searches, Higgs parameter measurements, and neutrino oscillation studies. Additionally, we provide a convenient ML tool that builds confidence intervals via a Neyman construction reliably and efficiently, with computational gains over traditional MC methods that are expected to scale with the dimensionality of the parameter space. The improved power of the FTS, enabled by the rapid building of the confidence intervals, can enhance the scientific discovery power of flagship measurements. Finally, our proposed method, which is here illustrated for histogram-based analyses, extends to unbinned, high-dimensional analyses with analytic expressions for the likelihood [Aalbers:2020iej] as well to as any other parametric [ATLAS:2018hxb] or nonparametric likelihood estimators [izbicki2014highdimensionaldensityratioestimation, Cranmer:2015bka, doi:10.1073/pnas.1912789117, ATLAS:2025clx, Ghosh:2025fma, dalmasso_likelihood-free_2024].
Acknowledgements.
This project was initiated through discussions at the PHYSTAT workshops in Munich and London in 2024, and we thank Mikael Kuusela, David Rousseau, Ibles Olcina, Callum Wilkinson, and Lukas Heinrich for fruitful conversations. We thank Scott Kravitz for guidance on setting up the LZ-inspired dataset and pointing us to relevant resources. We also thank Scott Kravitz and Ben Nachman for feedback on the manuscript. AG and DW were supported by the DOE Office of Science. JC and ABL were supported in part by NSF DMS-2053804. RI is grateful for the financial support of FAPESP (grant 2023/07068-1) and CNPq (grants 422705/2021-7 and 305065/2023-8). This work used Bridges-2 at PSC through allocation MTH250002 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program [10.1145/3569951.3597559], which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
Appendix A Experimental setup
Over its run, an experiment measures events generated via signal and background physics processes. Large-scale simulations of these processes yield template histograms that define the expected counts of signal and background events as functions of relevant observables. We denote these histograms ${\bm{\alpha}}(L)$ and ${\bm{\beta}}(L)$ for the signal and background distributions, respectively, where $L$ is the luminosity (a measure of total count) of the experiment. We model the number of events that fall into each histogram bin as a Poisson distribution. That is, the likelihood of the number of events $\mathcal{D}=(N_{1},N_{2},\ldots,N_{d})$ in $d$ bins has the form
$p(\mathcal{D}|\mu,{\bm{\alpha}},{\bm{\beta}})=\prod_{i=1}^{d}f(N_{i};\mu\alpha_{i}+\beta_{i}),$ (4)
where $\mu>0$ represents the signal strength, and $f(N;\lambda)$ denotes the probability mass function for a Poisson distribution with rate $\lambda$. We refer to each realization of $\mathcal{D}$ as a pseudo-experiment (PE).
The above “counting experiment” model convention is widely used for making likelihood-free inference tractable in cases where rare or exotic events are mixed with a relatively high volume of background events [cowan1998statistical, casadei_reference_2012]. In this Letter, we follow the standard convention of histogram analyses for such counting experiments.
A.1 Implementation of test statistics
We evaluate our test statistics on a grid of null parameters $\mu_{0}$ from $\Theta=[0,10]$ in our main results for the Higgs experiment (Section III.1), and from $[0,40]$ for the LZ-inspired experiment (Section III.2). The grid is always of a resolution of at least 1 point per 0.04 units in $\mu$. In Figures 1 and 2, we display results only on $[0,4]$ for Higgs and $[0,15]$ for LZ-inspired for the clear comparison of details.
For computing the LRS (Eq. 1), we replace $\Theta$ in the denominator with an extended parameter space,
$\widetilde{\Theta}=[\min(\Theta)-1,\max(\Theta)+1].$ (5)
This adjustment mitigates the effect of the boundary on the distribution of the LRS itself, independent of the method used for critical value estimation. The boundary effect arises because Wilks’ theorem, which guarantees that the critical values remain constant across the parameter space for LRS and large sample sizes (as in Figure 1 top), only holds for interior points [Algeri:2019lah, Herold:2024enb]. To ensure that the likelihood is well defined on $\mu<0$, we amend Eq. 4 to express
[…]
with rate parameter ensured to be positive. The supremum is computed using scipy’s implementation of the bounded Brent routine [2020SciPy-NMeth].
A.2 LZ-inspired dataset
We construct a dataset with signal and background shapes visually similar to those in Fig. 3 of Ref. [LZ:2024zvo]. Template histograms are generated using the nestpy Python interface (v2.0.4) to the NEST simulator (v2.4.0). The Lux_Run03 detector binding is adapted for this study, with the default parameter values used to determine the background electronic recoil (ER) template histogram and with set values of $g_{1}=0.092$ and $g_{1}^{gas}=0.076$ for the signal nuclear recoil (NR) histogram. These values were chosen to visually match the 40 GeV WIMP signal distribution in Ref. [LZ:2024zvo]. The resulting histograms are shown in Fig. 4. more detailed reproduction of the LZ setup or the use of official LZ simulation samples is beyond the scope of this Letter.
Figure 4: Template 2D histograms for the LZ-inspired dataset, showing the signal (nuclear recoil, NR; left) and background (electronic recoil, ER; right). Each histogram has 10 bins along both axes.
Appendix B Quantile regression
[…]
Data: test statistic $T(D;\mu)$; calibration data $\mathcal{T}=\{(\mu_{1},D_{1}),\ldots,(\mu_{B},D_{B})\}$; quantile regression estimator; level $\alpha\in(0,1)$
Result: estimated critical values $\widehat{C}_{\mu_{0}}$ for all $\mu_{0}\in\Theta$
Set $\widetilde{\mathcal{T}}_{\rm cal}\leftarrow\emptyset$;
for $i\in\{1,\ldots,B\}$ do
   Compute test statistic $t_{i}\leftarrow T(D_{i};\mu_{i})$;
 $\widetilde{\mathcal{T}}_{\rm cal}\leftarrow\widetilde{\mathcal{T}}_{\rm cal}\cup\{(\mu_{i},t_{i})\}$;
Use $\widetilde{\mathcal{T}}_{\rm cal}$ to learn the conditional quantile function $\widehat{C}_{\mu}=\widehat{F}_{T|\mu}^{-1}(\alpha|\mu)$ via quantile regression of $T$ on $\mu$;
return $\widehat{C}_{\mu_{0}}$
Algorithm 1 Estimate critical values $C_{\mu_{0}}$ for a level-$\alpha$ test
Appendix C Upper and lower limits for LZ-inspired experiment
For the LZ-inspired search, it is interesting to study the upper limits when there is no signal (i.e. $\mu^{*}=0$) and lower limits when there is a signal (i.e. $\mu^{*}\neq 0$). Figure 5 compares the upper and lower bounds for these respective scenarios for LRS and FTS. These results for the no-signal scenario are summarized in Table 2 and results for a scenario where the signal does exist with $\mu^{*}=10$ is summarized in Table 3. We find that when there is no signal, FTS is able to set tighter upper bounds, and when there is signal, FTS is able to set higher lower bounds compared to LRS.
$\mu^{*}$ LRS FTS, wide FTS, narrow
0.0 4.97 (12.59) 4.25 (10.98) 2.73 (11.14)
Table 2: Median upper bound for LZ-inspired experiment (in units of $\mu$) when there is no signal, $\mu^{*}=0$, at confidence level 68% (95%). The best result (the lowest median upper bound) is bold-faced.
$\mu^{*}$ LRS FTS, wide FTS, narrow
10.0 0.0 (0.0) 2.08 (0.0) 1.60 (0.0)
Table 3: Median lower bound for LZ-inspired experiment (in units of $\mu$) when the signal strength $\mu^{*}=10.0$, at confidence level 68% (95%). The best result (the highest median lower bound) is bold-faced.
Figure 5: Distributions of the upper and the lower bounds of 68% confidence intervals for the LZ-inspired study, shown as box plots — the median values are represented by circles (or stars) with the 25th and 75th percentiles of each distribution connected with a line. LRS (orange) is compared to FTS with wide focus (blue) over pseudo-experiments across different values of $\mu^{*}$ (x-axis). Both the median upper bound and the median lower bound are closer to the bisector for FTS than LRS, which is consistent with tighter parameter constraints.